two numbers differ by 5. if their sum is 19,then find the two numbers.
Answer by stanbon(75887)

two numbers differ by 5. if their sum is 19,then find the two numbers.
----
Equations:
x-y = 5
x+y = 19
----
Add and solve for "x":
2x = 24
x = 12
---
Solve for "y":
x + y = 19
12 + y = 19
y = 7
----
Solution: (12,7)
=====================
cheers,
Stan H.
==========